# Parallel Batch-Dynamic Maximal Matching with Constant Work per Update

### v1

@@ 3. Random Maximal Hyperedge Matching
In our dynamic algorithm, we must find a maximal matching on a subgraph in two situations: when inserting edges and during random settles. When inserting edges, any fast maximal matching would suffice. However, for the random settles, we need, among other properties, for the matched edges to be picked with sufficient randomness to protect them from the oblivious adversary. For brevity we use the same greedy algorithm for both cases. Sequential (static) greedy maximal matching randomly permutes the edges of a graph, then passes through them once, selecting an edge for the matching if all its endpoints are free (see pseudocode below). The initial permutation, which is chosen uniformly at random, can be generated in linear work and log 0pt (Gil et al., 1991). Parallel greedy maximal matching returns the same matching as the sequential greedy algorithm but processes many edges at once. Consider the dependence graph on the edges, where an edge $e$ depends on $e^{\prime}$ if $e$ is incident to $e^{\prime}$ and has lower priority than $e^{\prime}$. We call an edge a root of the dependence graph if it depends on no edge. In each round of the parallel algorithm, we add all roots to the matching, remove the matched edges and their neighbors, and update the root set. BFS provides a work efficient algorithm (for $r=2$) for matching in this way (Blelloch et al., 2012). In BFS they show that the dependence graph has 0pt $O(\log^{2}n)$ (Blelloch et al., 2012); this was improved to $O(\log n)$ by FN (Fischer and Noever, 2019).
However, a direct translation of BFS (Blelloch et al., 2012) gives $O(mr^{2})$ work for hypergraphs. We modify the method by which the root set is updated to yield $O(m^{\prime})$ work, where $m^{\prime}$ is the total cardinality of the hypergraph. We keep track of the following data structures. For each vertex $v$, neighbors$(v)$ gives a linked list of incident edges, edges$(v)$ gives an array of edges sorted by priority, and top$(v)$ gives the index of the highest priority remaining edge in edges$(v)$. We do not remove edges from edges$(v)$; instead, we increase top$(v)$ to indicate that prior edges are deleted. Each edge $e$ has pointers to each of its instances (up to $r$ each) in neighbors$(v)$ and edges$(v)$, and an atomic counter (counter), initially set to 0, indicating on how many vertices it is the highest priority remaining edge. We set done$(e)$ to be true if $e$ is being deleted this round, and false otherwise. Note that an edge is a root iff it is the highest priority remaining edge with respect to all of its vertices. When edges(v)[top$(v)$] is set to be deleted, we call updateTop$(v)$, which finds $v$’s new highest remaining edge $e^{\prime}$ and increments counter$(e^{\prime})$. If counter$(e^{\prime})$ equals the rank of $e^{\prime}$, we add it to the new root set. We repeat on the new root set and remaining graph. The overall algorithm is written in the function parallelGreedyMatch.
@@
$E^{\prime}\leftarrow E\mbox{ radix sorted by }\pi$
@@
for edges $\mbox{edges }e^{\prime}\mbox{ incident to }e$ :
@@
--Effect: Corrects top$(v)$ if needed
--Output: Returns edge to add to frontier in needed
@@
$D^{\prime}\leftarrow\mbox{groupBy}(\{(n,w):n\in N_{G}(w)\cup\{w\},w\in W\})$
$\mbox{minIn}\leftarrow\{(\arg\min_{e\in F}\pi(e)),n):(n,F)\in D^{\prime}\})$
$X\leftarrow X\cup\mbox{groupBy}(\mbox{minIn})$
$\mbox{finished}\leftarrow W\cup N_{G}(W)$
@@
$V_{f}\leftarrow\bigcup_{e\in\mbox{finished}}\{v|v\in e\}$
$W\leftarrow\{\mbox{updateTop}(v):v\in V_{f},\mbox{updateTop}(v)\neq\bot\}$
$E\leftarrow E\setminus\mbox{finished}$
@@
Figure 1. Sequential and parallel greedy maximal matching.
Define the sample space $S_{e}$ of a matched edge $e$ to be a set of unmatched edges we are assigning to $e$ [^3] The sample space corresponds to the temporarily deleted edges from Assadi and Solomon’s algorithm (Assadi and Solomon, 2021). The maximal matching we return with parallelGreedyMatch is augmented with sample spaces. For our dynamic algorithm to be efficient, we need for the sample spaces to have certain properties. To cover the cost of matching, we need all edges to be assigned to a sample space. We need an edge to belong to at most one sample space, so that its deletion only counts against the life of one epoch. Thus, we need the sample spaces to be disjoint. To ensure that edges are not reused in other sample spaces, we implicitly remove them from the graph until the matched edge is deleted. Thus we need for sampled edges to intersect their chosen matched edge, so that removing them does not violate maximality in future time steps. We formalize these desiderata in the following lemma.
@@
The function parallelGreedyMatch has the following properties.
  1. (1)
    $\{S_{e}:(e,S_{e})\in X\}$ partitions $E$.
  2. (2)
    For $(e,S_{e})\in X$, for all $e^{\prime}\in S_{e}$, $e\cap e^{\prime}\neq\emptyset$.
  3. (3)
    $\{e:(e,S_{e})\in X\}$ is a maximal matching on $G$.
@@
We can analyze sequentialGreedyMatch instead of parallelGreedyMatch because, as shown by BFS, they process edges in the same order (Blelloch et al., 2012).
  1. (1)
    Note that if $e^{\prime}$ is added to $S_{e}$ in time step $t$, $\mbox{free}(e^{\prime})=true$ before time step $t$ and $\mbox{free}(e^{\prime})=\mbox{false}$ after time step $t$. Thus the $S_{e}$ are disjoint.
  2. (2)
    Observe that if $\mbox{free}(e^{\prime})=\mbox{false}$, then $e^{\prime}\in S_{e}$ for some $e$. Consider the state of $\mbox{free}(e^{\prime})$ at time step $t$. If $\mbox{free}(e^{\prime})$ is true, $e^{\prime}$ is added to the matching, so $e^{\prime}\in S_{e^{\prime}}$. Otherwise, $e^{\prime}$ is already in some sample space $S_{e}$. Thus every edge belongs to a sample space. Thus the sample spaces partition $E$.
  3. (3)
    Note that we only add edges to the sample space that are incident on the matched edge.
Lemma 3.2.
All of the calls to updateTop cost, in aggregate, $O(m^{\prime})$ work, where $m^{\prime}$ is the total cardinality of the edges. Each call individually is $O(\log m)$ 0pt.
Proof.
Each call to updateTop$(v)$ slides $v$’s pointer along its edge list to the right, to the new highest priority remaining edge. Using the findNext operation, a call that slides a pointer $d$ distance costs $O(d)$. Because there are $m^{\prime}$ total edges across the lists, the pointers slide a total of $O(m^{\prime})$ distance, for $O(m^{\prime})$ work. Since the edge lists have at most $m$ members, each findNext call takes $O(\log m)$ 0pt, so updateTop has $O(\log m)$ 0pt. ∎
@@
Because the edges are given random priorities, bucket sorting takes $O(m)$ work in expectation and $O(\log m)$ 0pt whp (Blelloch et al., 2012; Cormen et al., 2001). Each edge $e$ is deleted once and costs $|e|$ work outside of updateTop, for $O(m^{\prime})$ work. In aggregate, internal calls to groupBy cost no more than $O(m)$ in expectation. Since updateTop in aggregate costs $O(m^{\prime})$, we have $O(m^{\prime})$ work total. Since updateTop and filter have $O(\log m)$ 0pt, each round of GreedyMatch is $O(\log m)$ 0pt. By Fischer and Noever there are $O(\log m)$ rounds whp (Fischer and Noever, 2019), for $O(\log^{2}m)$ 0pt total. ∎
@@ 3.1. Random selection in Greedy Maximal Matching
This is not true of parallelGreedyMatch, which chooses a matched edge and its sample space together, in the same random process. Note that we do not tell the adversary the sample space partitioning of a graph. But if we did tell the adversary the sample space of a matched edge, because the adversary knows the structure of the graph, the adversary would learn something about which edge was likely matched. For example, consider the path on 4 vertices $(1,2),(2,3),(3,4)$. Suppose that parallelGreedyMatch is run and the sample space of one of the matched edges is $\{(1,2),(2,3),(3,4)\}$. Since a matched edge is incident on all its sample space it follows that $(2,3)$ was matched. Thus, even knowing the size of the sample space reveals information about which edge was likely matched. We still can demonstrate that parallelGreedyMatch has the necessary randomness to protect its matched edges from the adversary, but we must approach the analysis differently.
To each matched edge $e\in E$, we assign price $|S_{e}|$; unmatched edges are given price 0. Suppose that the user deletes one edge $d_{t}$ per time step $i$ until all of the edges that participate in parallelGreedyMatch are deleted. If $p(d_{t})$, the matched edge whose sample space $d_{t}$ belongs to, has not yet been deleted (or if $d_{t}=p(d_{t})$), we call the delete “early.” Otherwise, we call the user delete “late.” When the user deletes an unmatched edge $e^{\prime}$, we pay 1 and subtract 1 from $p(e^{\prime})$’s price [^4] We only reduce $p(e^{\prime})$’s price if $p(e^{\prime})$ is still present in the graph (i.e., if $e^{\prime}$ is an early delete).. When the user visits a matched edge $e$, we pay the price of $e$. Let $\Phi(e)$ be the total price we pay when visiting edge $e$.
Remark: Consider a user delete of edge $d_{t}$. Suppose that $d_{t}$ is a late delete. Since $d_{t}$’s matched edge is gone, $d_{t}$ is unmatched. Thus while visiting $d_{t}$ we will pay 1 unit of price. Thus $\mathbb{E}[\Phi(d_{t})]=1$. Note that we do not need this fact, as late deletes do not appear in our analysis later.
@@
If $d_{t}$ is an early delete, then $\mathbb{E}[\Phi(d_{t})]\leq 2$.
@@
We consider sequentialGreedyMatch because it processes edges in the same order as the parallel version (Blelloch et al., 2012). Let UserRemain be the remaining incident edges on $d_{t}$ at (the beginning of) time $t$. The other edges that were incident on $d_{t}$ at time 0 have been deleted by the user. Let $\Pi$ be the set of permutations (that we would input into sequentialGreedyMatch) such that $d_{t}$ is an early delete. Then $\Pi$ is the set of permutations such that $p(d_{t})\in\mbox{UserRemain}$. Let $r$ be the time step in sequentialGreedyMatch at which $d_{t}$ is marked to be not free (the time step where $d_{t}$ is effectively removed from the graph). [^5] Note that we have two distinct uses of time: the time in the user delete sequence, and the time in the run of sequentialGreedyMatch. Note that the sequentialGreedyMatch time happens completely before the user delete sequence time. Note that these times are separate from each other: if $t$ is large, it is possible that $r$ is large or small. Let PermRemain be the set of edges incident on $d_{t}$ at step $r$. Group $\Pi$ into equivalence classes based on, for $\pi\in\Pi$, the value of $[r,(\pi_{i})_{i=1}^{r-1}]$. Note that PermRemain is determined by $[r,(\pi_{i})_{i=1}^{r-1}]$. Let $\pi\in\Pi$, and let $B$ be the equivalence class for which $\pi\in B$.
We consider the possible values of $\pi^{-1}(r)$. Since $d_{t}$ must delete in step $r$, $\pi^{-1}(r)$ must be in PermRemain. Note that $\pi^{-1}(r)=p(d_{t})\in\mbox{UserRemain}$. Since there are no other restrictions, we have that $\pi^{-1}(r)\in\mbox{PermRemain}\cap\mbox{UserRemain}$. Thus $|B|=|\mbox{PermRemain}\cap\mbox{UserRemain}|(m-r)!$. Next, we consider the value of $\Phi(d_{t})$ if $d_{t}$ is matched. At user time 0, $d_{t}$ had price —PermRemain—. However, each unmatched edge in $d_{t}$’s sample space that the user visits reduces $d_{t}$’s price by 1. Thus, the price of $d_{t}$ is $|\mbox{PermRemain}|-|\mbox{PermRemain}\cap\overline{\mbox{UserRemain}}|=|\mbox{PermRemain}\cap\mbox{UserRemain}|$. There are $(m-r)!$ permutations in $B$ where $d_{t}$ matches at time step $r$.
Drawing a permutation uniformly at random from $B$, we get $\mathbb{E}[\Phi(d_{t})]=|\mbox{PermRemain}\cap\mbox{UserRemain}|(1/|\mbox{PermRemain}\cap\mbox{UserRemain}|)+(1-1/|\mbox{PermRemain}\cap\mbox{UserRemain}|)1\leq 2$. Since $B$ was an arbitrary equivalence class we conclude that $\mathbb{E}[\phi(d_{t})]=\mathbb{E}[\phi(d_{t})|\pi\in B]\leq 2$. ∎
Note that Lemma 3.4 is true for all times $t$; this means that, regardless of the deletes the user chooses, the next delete has constant expected price. Even though the price of later deletes depends on the price of earlier deletes, the expectation remains no more than 2. We also note that, when a graph is fully deleted, the sum of the prices of early deletes is the number of edges in the graph. Let $\Phi^{\prime}(d_{t})=\begin{cases}\Phi(d_{t})&d_{t}\textrm{ is early}\\ 0&\textrm{otherwise}\end{cases}$.
@@
The early deletes on the sample space of deleted matched edge $e$ contribute $|S_{e}|$ total price. Furthermore, $\sum_{i=1}^{m}\Phi^{\prime}(d_{t})=m$.
@@
Observe that if a matched edge $e$ has price $|S_{e}|-j$ on deletion, then $j$ unmatched edges in $S_{e}$ were deleted, and the updates deleting these $j$ unmatched edges paid $j$ price. Thus the early deletes on the sample space of $e$ contribute $|S_{e}|$ total price. Thus the price of all early deletes is $m$. ∎
isHeavy$(e)$ : return $|C(e)|\geq 4r^{2}2^{l(e)}$
@@
addCrossEdge$(e)$ :
type$(e)\leftarrow$cross
$v_{max}\leftarrow argmax_{v\in V(e)}l(p(e))$
@@
removeCrossEdge$(e)$ :
$C(p(e))\leftarrow C(p(e))\setminus\{e\}$
parfor $v\in V(e)$ : delete$(P(v,l(p(e))),e)$
type$(e)\leftarrow$unsettled
@@
parfor $e\in C$ : removeCrossEdge$(e)$
parfor $e\in C$ : addCrossEdge$(e)$
@@
deleteMatchedEdges$(E)$ :
parfor $e\in\bigcup_{m\in E}S(m)$ : addCrossEdge$(e)$
@@
insertEdges$(\bigcup_{m\in\mbox{light}}\mbox{removeMatch}(m))$
@@
[…]
parfor $e\in Y$ : addMatch$(e,\{e\})$
parfor $e\in E\setminus Y$ : addCrossEdge$(e)$
@@
if $\mbox{type}(e)=\mbox{cross}$ : removeCrossEdge$(e)$
@@
parfor $m\in\mbox{matched}$ : delete$(S(m),m)$
@@
insertEdges$(E^{\prime})$
Figure 2. Parallel batch-dynamic algorithm for dynamic matching with $O(r^{3})$ expected amortized cost per edge update.

### v2

@@ 3. Random Maximal Hyperedge Matching
In our dynamic algorithm, we must find a maximal matching on a subgraph in two situations: when inserting edges and during random settles. When inserting edges, any fast maximal matching would suffice. However, for the random settles, we need, among other properties, for the matched edges to be picked with sufficient randomness to protect them from the oblivious adversary. For brevity we use the same algorithm for both cases, and in particular we develop an efficient parallel implementation of the static sequential random greedy algorithm for this purpose.
The sequential random greedy maximal matching algorithm first randomly permutes the edges of a graph, and then passes through them once in the random order, selecting an edge if it has not been deleted, and if it is selected deletes all incident edges. We refer to the order in the permutation as the priority (highest first). The output of this algorithm is known as the the lexiographically first maximal matching (Cook, 1985). For our purposes, it is important to keep track of the edges that are deleted when a matched edge is selected. For a matched edge $e$ we refer to these edges as the sample space $S_{e}$ of $e$. The sample space includes $e$ itself. Note that the sample spaces of the matched edges form a partitioning of the original edges since every edge is deleted exactly once. Figure 1 gives pseudocode for the algorithm.
@@
$E$ $^{\prime}\leftarrow E\mbox{ sorted by }\pi$
@@
for $\mbox{edges }e^{\prime}\mbox{ incident to }e$ :
@@
Figure 1. Sequential random greedy maximal matching.
Greedy maximal matching can be implemented to run in parallel, returning the identical output. We call an edge $e$ a root of a graph if all remaining edges incident to $e$ have later priority than $e$. In each round of the parallel algorithm, we add all roots ($W$) to the matching, remove the matched edges along with their incident edges, and update the root set. Since edges in the root set can have mutually neighboring edges, to mimic the sample spaces of the sequential algorithm, we assign such a contested edge to the sample space of the highest priority incident root. Blelloch, Fineman and Shun (BFS) show that this procedure takes $O(\log^{2}n)$ rounds whp (Blelloch et al., 2012); and the analysis was improved to $O(\log n)$ rounds whp by Fischer and Noever (FN) (Fischer and Noever, 2019). Both of these results reduce matching to finding a maximally independent set of edges, where independence is with respect to edge incidence. For $r=2$, BFS also provide a work efficient (linear work) parallel algorithm for this approach that runs in polylogarithmic depth. A direct translation of this algorithm to hypergraphs, however, is not work efficient.
Here we develop a work-efficient maximal matching algorithm on hypergraphs by modifying the method by which the root set is updated to yield $O(m^{\prime})$ work, where $m^{\prime}$ is the total cardinality of the hypergraph. Towards this end we keep track of the following data structures. For each vertex $v$, we store the edges incident on $v$ in two data structures. We maintain the remaining edges adjacent to $v$, $N(v)$, as an unordered set; when an edge incident to $v$ is removed from the graph, we remove this edge from $N(v)$. We also maintain edges(v)(v), an array of edges sorted by priority, and top(v)(v), the index of the highest priority remaining edge in edges(v)(v). We do not remove edges from edges(v)(v); instead, we increase top(v)(v) to indicate that prior edges are deleted. Each edge $e$ has an atomic counter (counter), initially set to zero, indicating on how many vertices it is the highest priority remaining edge. We set done(e)(e) to be true if $e$ is being deleted this round, and false otherwise.
[…]
@@
--Effect: Corrects top(v)(v) if needed
--Output: Returns edge to add to frontier if needed
@@
$D\leftarrow\mbox{groupBy}(\{(e,w):w\in W,e\in N(V(w))\})$
$X^{\prime}\leftarrow\mbox{groupBy}(\{(\arg\min_{e\in F}\pi(e)),e):(e,F)\in D\})$
$X\leftarrow X\cup X^{\prime}$
$\mbox{finished}\leftarrow W\cup N(V(W))$
@@
$V_{f}\leftarrow\bigcup_{e\in\mbox{finished}}V(e)$
$W^{\prime}\leftarrow\{\mbox{updateTop}(v):v\in V_{f}\}$
$W\leftarrow\{e\in W^{\prime}\;|\;e\neq\bot\}$
parfor $e\in\mbox{finished},v\in V(e)$ : $\mbox{delete}(N(v),e)$
@@
Figure 2. Parallel random greedy maximal matching. In the code, $V(e)$ indicates the vertices incident on $e$, and for a set of vertices $V$, $N(V)$ indicates the union of all edges incident on at least one vertex in $V$.
We now analyze the cost of parallelGreedyMatch.
@@
All of the calls to updateTop cost, in aggregate, $O(m^{\prime})$ work, where $m^{\prime}$ is the total cardinality of the edges and $m$ is the total number of edges. Each call individually is $O(\log m)$ 0pt.
@@
Each call to updateTop(v)(v) slides $v$’s pointer along its edge list to the right, to the new highest priority remaining edge. Using the findNext operation, a call that slides a pointer $d$ distance costs $O(d)$. Because there are $m^{\prime}$ total edges across the lists, the pointers slide a total of $O(m^{\prime})$ distance, for $O(m^{\prime})$ work. Since the edge lists have at most $m$ members, each findNext call takes $O(\log m)$ 0pt, so updateTop has $O(\log m)$ 0pt. ∎
@@
Generating the permutation for the edges takes linear work. Because the edges are given random priorities, bucket sorting takes $O(m)$ work in expectation and $O(\log m)$ 0pt whp (Blelloch et al., 2012)(Cormen et al., 2009, p. 200-204). Each edge $e$ is deleted once and costs $|e|$ work outside of updateTop, for $O(m^{\prime})$ work. In aggregate, internal calls to groupBy cost no more than $O(m)$ in expectation. Since updateTop in aggregate costs $O(m^{\prime})$, we have $O(m^{\prime})$ work total. Since updateTop and filter have $O(\log m)$ 0pt, each round of GreedyMatch is $O(\log m)$ 0pt. By FN there are $O(\log m)$ rounds whp (Fischer and Noever, 2019), for $O(\log^{2}m)$ 0pt total. ∎
@@ 3.1. Random Selection in Greedy Maximal Matching
This is not true of parallelGreedyMatch, which chooses a matched edge and its sample space together, in the same random process. Note that we do not tell the adversary the sample space partitioning of a graph. But if we did tell the adversary the sample space of a matched edge, because the adversary knows the structure of the graph, the adversary would learn something about which edge was likely matched. For example, consider the path on 4 vertices $(1,2),(2,3),(3,4)$. Suppose that parallelGreedyMatch is run and the sample space of one of the matched edges is $\{(1,2),(2,3),(3,4)\}$. Since a matched edge is incident on all its sample space it follows that $(2,3)$ was matched. Thus, even knowing the size of the sample space reveals information about which edge was likely matched. We still can demonstrate that parallelGreedyMatch has the necessary randomness to protect its matched edges from the adversary, but we must approach the analysis differently. Note that the analysis presented in this section is independent of the dynamic matching that follows. In particular, in this setting, induced and natural epochs do not exist. We use the results from this section as a black box in Section 5.3 to show bounds on the dynamic algorithm.
We start by presenting an alternative simple proof for the case where the sample space is chosen deterministically, and then extend to the general case. [^5] This proof seems significantly simpler than the proofs in (Solomon, 2016; Assadi and Solomon, 2021). To each matched edge $e\in E$, we assign price $|S_{e}|$, which has to be paid for. Unmatched edges are given price 0. In the dynamic matching setting, after an element is chosen to be matched from the sample space, the user deletes one edge $d_{t}$ per time step $t$ until all of the edges of the sample space are deleted. [^6] The amortized analysis assumes the graph ends up empty. The user does these deletes in an arbitrary and adversarial order, but since the user is oblivious, she does not know which edge in the sample space was matched. Let $p(e)$ indicate the matched edge for which $e$ is in its sample space ($S_{p(e)}$), allowing us later to consider multiple matched edges and corresponding sample spaces.
We say that when the user deletes an edge $d_{t}$ after $p(d_{t})$ was deleted, that the deletion of $d_{t}$ is late, and otherwise it is early. Note that if $d_{t}=p(d_{t})$, then it is early. We need to pay the price for a matched edge with the user deletes on its sample space. We cannot, however, use late deletes to pay the price since after the match is deleted a new incident edge is matched and the future delete of an edge might have to pay for that one. We therefore charge all the price of a match to early deletes on edges in its sample space, and will argue that with an expected payment of at most 2 units per delete we can cover the full price.
We use the following charging scheme. When the user deletes an early unmatched edge $e$, it pays 1 unit to subtract 1 unit from $p(e)$’s price. When the user deletes a matched edge $e$, which must be early, it pays the remaining price of ee—i.e., the original price minus the number of edges in its sample space that have been visited. If $e$ is late it pays nothing. Note that in this scheme once all edges have been deleted the payment of early deletions fully covers the price. Let $\Phi(e)$ be the random variable for the payment made when deleting the edge, and $U(e)$ be the set of edges in the sample set of $e$ just before it is deleted. We are interested in $\mathbb{E}[\Phi(e)]$ for early deletions.
Proof.
(For deterministically chosen sample spaces, if $d_{t}$ is an early delete, then $\mathbb{E}[\Phi(d_{t})]<2$.) Given the two cases for costs for early deletes, we have that
$E[\Phi(e)]=p_{\mbox{\small unmatched}}\times 1+p_{\mbox{\small matched}}\times|U(e)|\;.$
However $p_{\mbox{\small matched}}$ is exactly $1/|U(e)|$, since each remaining edge is equally likely to have been picked from the sample space. Hence we have $E[\Phi(e)]<1\times 1+\frac{1}{|U(e)|}\times|U(e)|=2$. ∎
Note that to extend this argument to our more complicated situation where the sample space is itself randomized, the first term remains the same since the probability is always upper bounded by 1. The second term, however, is more complicated since “each remaining edge is equally likely to have been picked from the sample space” is not well defined since the sample space depends on the random permutation in our matching. Intuitively, it still works because the possible permutations can be partitioned into equivalence classes each consisting of a fixed sample space, and within each equivalence class we can argue that every element in the sample space is equally likely. We formalize this idea in the following theorem and proof.
@@
For sample spaces created by sequentialGreedyMatch, if $d_{t}$ is an early delete, then $\mathbb{E}[\Phi(d_{t})]\leq 2$.
@@
Let $\Pi$ be the set of permutations (that we would input into sequentialGreedyMatch) such that $d_{t}$ is an early delete. Then $\Pi$ is the set of permutations such that $p(d_{t})\in U(d_{t})$ (henceforth we will use $U$ for $U(d_{t})$). Let $r$ be the time step in sequentialGreedyMatch at which $d_{t}$ is marked to be not free (the time step where $d_{t}$ is effectively removed from the graph). [^7] Note that we have two distinct uses of time: the time in the user delete sequence, and the time in the run of sequentialGreedyMatch. Note that the sequentialGreedyMatch time happens completely before the user delete sequence time. Note that these times are separate from each other: if $t$ is large, it is possible that $r$ is large or small. Let $P$ be the set of edges incident on $d_{t}$ at step $r$ (the remaining edges after permutation deletes). Group $\Pi$ into equivalence classes based on, for $\pi\in\Pi$, the value of $r,(\pi_{i})_{i=1}^{r-1}]$. Note that $P$ is determined by $r,(\pi_{i})_{i=1}^{r-1}]$. Let $\pi\in\Pi$, and let $B$ be the equivalence class containing $\pi$.
We consider the possible values of $\pi^{-1}(r)$. Since $d_{t}$ must delete in step $r$, $\pi^{-1}(r)$ must be in $P$. Note that $\pi^{-1}(r)=p(d_{t})\in U$. Since there are no other restrictions, we have that $\pi^{-1}(r)\in P\cap U$. Thus $|B|=|P\cap U|(m-r)!$. Next, we consider the value of $\Phi(d_{t})$ if $d_{t}$ is matched. At user time 0, $d_{t}$ had price $|P|$. However, each unmatched edge in $d_{t}$’s sample space that the user visits reduces $d_{t}$’s price by 1. Thus, the price of $d_{t}$ is $|P|-|P\cap\overline{U}|=|P\cap U|$. There are (mr)!(m-r)! permutations in $B$ where $d_{t}$ matches at time step $r$.
Drawing a permutation uniformly at random from $B$, we get
$\mathbb{E}[\Phi(d_{t})]=(1-\frac{1}{|P\cap U|})1+\frac{1}{|P\cap U|}|P\cap U|\leq 2.$
Since $B$ was an arbitrary equivalence class we can conclude that $\mathbb{E}[\phi(d_{t})]\leq\max_{B}\mathbb{E}[\phi(d_{t})|\pi\in B]\leq 2$. ∎
Note that Lemma 3.3 is true for all times $t$; this means that, regardless of the deletes the user chooses, the next delete has constant expected payment. Even though the payment given by later deletes depends on the payment given by earlier deletes, the expectation remains no more than 2. We also note that, when a graph is fully deleted, the sum of the payments of early deletes is the number of edges in the graph.
Let $\Phi^{\prime}(d_{t})=\begin{cases}\Phi(d_{t})&d_{t}\textrm{ is early}\\ 0&\textrm{otherwise}\end{cases}$.
@@
The early deletes on the sample space of deleted matched edge $e$ contribute $|S_{e}|$ total payment. Furthermore, $\sum_{i=1}^{m}\Phi^{\prime}(d_{t})=m$.
@@
Observe that if a matched edge $e$ has price $|S_{e}|-j$ on deletion, then $j$ unmatched edges in $S_{e}$ were deleted, and the updates deleting these $j$ unmatched edges paid $j$. Thus the early deletes on the sample space of $e$ contribute $|S_{e}|$ total payment. Thus, all of these early deletes pay $m$ in total. ∎
isHeavy(e)(e) : return $|C(e)|\geq 4r^{2}2^{l(e)}$
@@
addCrossEdge(e)(e) :
[…]
$v_{max}\leftarrow argmax_{v\in V(e)}l(p(v))$
@@
removeCrossEdge(e)(e) :
$\mbox{delete}(C(p(e)),e)$
parfor $v\in V(e)$ : delete(P(v,l(p(e))),e)(P(v,l(p(e))),e)
[…]
@@
parfor $e\in C$ : removeCrossEdge(e)(e)
parfor $e\in C$ : addCrossEdge(e)(e)
@@
deleteMatchedEdges(E)(E) :
parfor $e\in\bigcup_{m\in E}S(m)$ : addCrossEdge(e)(e)
@@
[…]
@@
$Y\leftarrow\{e:(e,\_)\in\mbox{parallelGreedyMatch(G')}\}$
parfor $e\in Y$ : addMatch(e,{e})($\{e\}$)
parfor $e\in E\setminus Y$ : addCrossEdge(e)(e)
@@
if $\mbox{type}(e)=\mbox{cross}$ : removeCrossEdge(e)(e)
@@
parfor $m\in\mbox{matched}$ : delete(S(m),m)(S(m),m)
@@
[…]
Figure 3. Parallel batch-dynamic algorithm for maximal matching with $O(r^{3})$ expected amortized cost per edge update.
@@
Figure 4. Flow chart representation of deleteEdges. After edges are removed from the graph, they are separated into matched and unmatched. The matched edges are passed into deleteMatchedEdges, where they are separated into light and heavy. The light matches (and their cross edges) are reinserted into the graph, while the heavy matches and their cross edges ($E^{\prime}$) are sent to randomSettle. A static maximal matching is run on $E^{\prime}$, and the matches are separated into satisfied (no further action needed), stolen, and bloated. We repeat the process by sending the stolen and bloated edges to deleteMatchedEdges. We repeat until we have sampled sufficiently many edges.